# Fast Polynomial Root-Finding and Approximation of the Eigenvalues of a Matrix

### v36

@@ 26.3 Acceleration with Fast Multipole Method (FMM)
The algorithms of [81, 59] deeply involve the coefficients of $p(x)$. Next we cover acceleration of Newton’s, Weierstrass’s, and Ehrlich’s iterations applied to black box polynomial root-finding. Given $3d$ complex values $w_{j}$, $z_{i}$, $v_{j}$ such that $z_{i}\neq v_{j}$ for all $i,j=1,\dots,d$, Fast Multipole Method (FMM) by Greengard and Rokhlin [50] and by Carrier, Greengard, and Rokhlin [28] computes the sums $\sum_{j=1}^{d}\frac{w_{j}}{x_{i}-z_{j}}$ as well as $\sum_{j=1,j\neq i}^{d}\frac{w_{j}}{z_{i}-z_{j}}$ for all $i$ by using $O(db)$ arithmetic operations, with the precision of order $b$, versus $2d^{2}$ arithmetic operations in the straightforward algorithm. The paper [121] proposes and extensively tests its acceleration to Superfast Multipole Method. The papers [105, 107] list the recent bibliography on FMM and related computations with Hierarchical Semiseparable (HSS) matrices. Relevant software libraries developed at the Max Planck Institute for Mathematics in the Sciences can be found in HLib,
@@
Next we incorporate FMM into Newton’s, Weierstrass’s, and Ehrlich’s iterations. First relate the roots of a polynomial $p(x)$ to the eigenvalues of a diagonal-plus-rank-one matrix
@@
Every iteration but the first one begins with some values $z_{j}$, $w_{j}$, and updated values $z_{j}^{\rm(new)}$ of (26.1), for $j=1,\dots,d$, and outputs updated values $w_{j}^{\rm(new)}=W(z_{j}^{\rm(new)})$ for all $j$. Then equations (26.3) and (26.5) define $N(z_{j}^{\rm(new)})$ and $E_{j}(z_{j}^{\rm(new)})$ for all $j$. To update the values $w_{j}$, apply [22, equation (17)]:
@@
(i) Use the following customary trick (see [32, Remark 3]): approximate the values $\ln(R_{i})$ for $i=1,\dots,d$ fast by applying FMM and then readily approximate $R_{i}=\exp(\ln(R_{i}))$ for all $i$. [32] proves numerical stability of the reduction to FMM of a similar albeit a little more involved rational expressions.
@@
By applying FMM fast approximate $\sigma_{i}$ for all $i$ and compute the values $\prod_{j\in\mathbb{J}^{\prime}}\frac{z_{i}^{\rm new}-z_{j}}{z_{i}^{\rm new}-z_{j}^{\rm new}}$ for all $i$ in at most $4\sum_{i=1}^{d}~{}|\mathbb{J}_{i}^{\prime}|-d$ arithmetic operations for $|\mathbb{J}^{\prime}_{i}|$ denoting the cardinality of the set $\mathbb{J}^{\prime}_{i}$. Compare the latter cost bound with order of $d^{2}$ arithmetic operations in the straightforward algorithm.
@@
Let us comment on the changes where one incorporates FMM into Newton’s, Weierstrass’s, and Ehrlich’s iterations implicitly applied to the wild factor $f(x)$ of $p(x)$. The same expression (3.3) defines Ehrlich’s iterations for both $f(x)$ and $p(x)$ by virtue of Thm. 13.1. Newton’s iterations to a wild factor one additionally subtract the sum $\sum_{j=1}^{d-w}\frac{1}{x-z_{j}}$ (cf. (3.3)) and can apply FMM to accelerate numerical computation of this sum. To apply Weierstrass’s iterations to a wild factor one performs additional division by $t(x)$ and decreases the degree of the divisor $l^{\prime}(x)$ from $d-1$ to $w-1$.
@@ 26.4 MPSolve and its competition with Newton’s iterations
Since its creation in 2000 the package MPSolve by Bini and Fiorentino [15], revised in 2014 by Bini and Robol [22], has been remaining user’s choice for approximation of all $d$ complex roots of a polynomial $p$. It implements Ehrlich’s iterations of 1967 [38], aka Aberth’s of 1973 [3], which empirically converge very fast globally – right from the start, under proper initialization, involving the coefficients of $p$. Such convergence has been consistently observed in decades-long worldwide applications of these iterations as well as in the case of some other root-finders by means of functional iterations, notably Weierstrass’s iterations [148], aka Durand and Kerner’s [31, 67]).
@@
Serious competition to MPSolve came in 2001 from the package EigenSolve by Fortune [43], but the later version of MPSolve [22] combines the benefits of both packages. An implementations of Newton’s iterations for approximation of all roots by Schleicher and R. Stoll in [143] competed with MPSolve in 2017, but then MPSolve has regained upper hand by means of incorporation of Fast Multipole Method (FMM), even though such implementation is simpler and can be more efficient for Newton’s than Ehrlich’s iterations (see the previous subsection). The competition seems to resume in 2024 (see [12, 130, 85]).
@@ 26.5.1 Introductory comments
Given coefficients of a polynomial $p$ of a degree $d$, we can readily compute its largest and smallest root radii $r_{1}$ and $r_{d}$, respectively (cf. [41, 52]). Next we compute a range $[\rho_{-},\rho_{+}]$ that brackets all root radii or, equivalently, compute an upper estimate for the largest root radius $r_{1}$ and a low estimate for the smallest root radius $r_{d}$ of $p(x)$.
We can apply the estimates of Sec. 13.4 to the reverse polynomial $p_{\rm rev}(x)$ if its coefficients are available, but next we study this problem for a black box polynomial. Its solution is required, e.g., in order to initialize algorithms for all $d$ roots by means of functional iterations.
Suppose that we have an algorithm that computes a disc $D(0,\rho_{-})$ containing no zeros of a black box polynomial $p(x)$; clearly, $\rho_{-}\leq r_{d}$. By applying this algorithm to the reverse polynomial $p_{\rm rev}(x)$ we will obtain the upper bound $1/\rho_{-,{\rm rev}}$ on $r_{1}$.
@@
Hereafter let $p(x)$ denote a black box polynomial.
@@ 26.5.3 Deterministic solution
We can deterministically compute a desired disc $D(0,\rho_{-})$ at a higher cost, based on a soft e/i test.
[…]
@@ 27 Fast approximation of the $d$ root radii with applications
The customary initialization recipes of Weierstrass"s and Ehrlich’s iterations involve the $d$ root radii at the origin and consequently require to use the coefficients of $p$. To initialize Ehrlich’s iterations MPSolve first applies Bini’s heuristic recipe of [10], reused in [15, 22], for the approximation of the $d$ root radii. Given such approximations MPSolve defines $k$ concentric annuli $A(c,\rho_{i},\rho_{i}^{\prime})$ and $k$ integers $m_{i}$ such that #$\#(A(c,\rho_{i},\rho_{i}^{\prime}))=m_{i}$ for $i=1,\dots,k$, $k\leq d$, and $\sum_{i=1}^{k}m_{i}=d$. Then MPSolve initializes Ehrlich’s iterations at $d$ points $z_{j}$, $j=1,2,\dots,d$, exactly $m_{i}$ of them being equally spaced on the circle $C(c,(\rho_{i}+\rho_{i}^{\prime})/2)$. This heuristic initialization rec;ipe turns out to be highly efficient.
The algorithm supporting the next theorem also approximates the $d$ root radii but is certified, deterministic, and faster than Bini’s. It was first outlined and briefly analyzed in [134] for fast approximation of a single root radius and then extended in [92, Sec. 4] (cf. also [93, Sec. 5]) to approximation of all root radii of a polynomial. As in [10, 15, 22] it uses all $d$ coefficients of $p$. A variation of this algorithm in [62] supports the following theorem.
@@
[62, Prop. 3]. Given the coefficients of a polynomial $p=p(x)$, one can approximate all the $d$ root radii $|x_{1}|,\dots,|x_{d}|$ within the relative error bound $4d$ at a Boolean cost in $O(d\log(||p||_{\infty}))$.
@@
Combine Thm. 27.1 with estimates for the Boolean cost of performing these iterations implicit in the proof of [134, Cor. 14.3] and obtain
@@
The algorithm supporting the corollary and said to be Alg. 27.1a computes at a low cost $d$ suspect annuli $A(0,r_{j,-},r_{j,+})$ for $j=1,\dots,d$ (some of them may pairwise overlap), whose union $\mathbb{U}$ contains all $d$ roots. The algorithm was the point of departure for amazing progress in [59], which extends the novel ingenious algorithm of [81] for multipoint polynomial evaluation.
@@
(ii) The search for real roots can be narrowed to the intersection $\mathbb{I}$ of $\mathbb{U}$ with the real axis, which consists of $2m$ segments of small length (see Fig. 11). [126] has tested a simple real root-finder that approximates the roots in $\mathbb{I}$ by initiating Newton’s iterations at the centers of the segments. According to these tests the algorithm is highly efficient for the approximation of all isolated real roots. One can extend Newton’s iterations at the centers of the segments by applying also our root-radii algorithm of Sec. 9. By complementing the root radii approach with some advanced novel techniques, the paper [62] has noticeably accelerated the real root-finder [70], which is currently user’s choice.
(iii) Initialization based on Alg. 27.1a should improve global convergence of Ehrlich’s, Weierstrass’s, and other functional iterations for root-finding as well as of Kirrinnis’s factorization algorithm of [68].
[…]

### v37

@@ 24.3 Acceleration with Fast Multipole Method (FMM)
The algorithms of [64, 44] deeply involve the coefficients of $p(x)$. Next we cover acceleration of Newton’s, Weierstrass’s, and Ehrlich’s iterations applied to black box polynomial root-finding. Given $3d$ complex values $w_{j}$, $z_{i}$, $v_{j}$ such that $z_{i}\neq v_{j}$ for all $i,j=1,\dots,d$, FMM computes the sums $\sum_{j=1}^{d}\frac{w_{j}}{x_{i}-z_{j}}$ as well as $\sum_{j=1,j\neq i}^{d}\frac{w_{j}}{z_{i}-z_{j}}$ for all $i$ by using $O(db)$ arithmetic operations, with the precision of order $b$, versus $2d^{2}$ arithmetic operations in the straightforward algorithm. The papers [86, 88] list the recent bibliography on FMM and related computations with Hierarchical Semiseparable (HSS) matrices. Relevant software libraries developed at the Max Planck Institute for Mathematics in the Sciences can be found in HLib,
@@
Next we incorporate FMM into Newton’s, Weierstrass’s, and Ehrlich’s iterations. First relate the roots of a polynomial $p(x)$ to the eigenvalues of a diagonal-plus-rank-one matrix [^31] Hereafter T denotes transposition of a vector.
@@
Every iteration but the first one begins with some values $z_{j}$, $w_{j}$ and the updated values $z_{j}^{\rm(new)}$ of (24.1), for $j=1,\dots,d$, and outputs the updated values $w_{j}^{\rm(new)}=W(z_{j}^{\rm(new)})$ for all $j$. Then equations (24.3) and (24.5) define $N(z_{j}^{\rm(new)})$ and $E_{j}(z_{j}^{\rm(new)})$ for all $j$. To update the values $w_{j}$, apply [15, equation (17)]:
@@
(i) Use the following customary trick (see [24, Remark 3]): approximate the values $\ln(R_{i})$ for $i=1,\dots,d$ fast by applying FMM and then readily approximate $R_{i}=\exp(\ln(R_{i}))$ for all $i$. [24] proves numerical stability of the reduction to FMM of a similar although a little more involved rational expressions.
@@
By applying FMM approximate $\sigma_{i}$ for all $i$ and compute the values $\prod_{j\in\mathbb{J}^{\prime}}\frac{z_{i}^{\rm new}-z_{j}}{z_{i}^{\rm new}-z_{j}^{\rm new}}$ for all $i$ in at most $4\sum_{i=1}^{d}~{}|\mathbb{J}_{i}^{\prime}|-d$ arithmetic operations for $|\mathbb{J}^{\prime}_{i}|$ denoting the cardinality of the set $\mathbb{J}^{\prime}_{i}$. Compare the latter cost bound with order of $d^{2}$ arithmetic operations in the straightforward algorithm.
@@
Let us comment on the changes where one incorporates FMM into Newton’s, Weierstrass’s, and Ehrlich’s iterations implicitly applied to the wild factor $f(x)$ of $p(x)$. The same expression (3.3) defines Ehrlich’s iterations for both $f(x)$ and $p(x)$ by virtue of Thm. 6.1. Applying Newton’s iterations to a wild factor one additionally subtract the sum $\sum_{j=1}^{d-w}\frac{1}{x-z_{j}}$ (cf. (3.3)) and can apply FMM to accelerate numerical computation of this sum. Applying Weierstrass’s iterations to a wild factor one performs additional division by $t(x)$ and decreases the degree of the divisor $l^{\prime}(x)$ from $d-1$ to $w-1$.
@@ 24.4 MPSolve and its competition with Newton’s iterations
Since its creation in 2000 the package MPSolve by Bini and Fiorentino [9], revised in 2014 by Bini and Robol [15], has been remaining user’s choice for approximation of all $d$ complex roots of a polynomial $p$. It implements Ehrlich’s iterations, which empirically converge very fast globally, right from the start, under proper initialization involving the coefficients of $p$. Such convergence has been consistently observed in decades-long worldwide applications of these iterations as well as in the case of some other root-finders by means of functional iterations, notably Weierstrass’s iterations.
@@
Serious competition to MPSolve came in 2001 from the package EigenSolve by Fortune [33], but the later version of MPSolve [15] combines the benefits of both packages. An implementations of Newton’s iterations for approximation of all roots by Schleicher and R. Stoll in [115] competed with MPSolve in 2017, but then MPSolve has regained upper hand by means of incorporation of FMM. As we showed in the previous subsection, however, such implementation is simpler and can be more efficient for Newton’s than Ehrlich’s iterations. The Ehrlich-Newton competition seems to resume in 2024 (see [7, 102, 68]).
@@ 24.5 Computation of a disc covering all zeros of a black box polynomial
Customary initialization of Ehrlich’s and Weierstrass’s iterations involves approximation of all root radii, for which we can $d$ times apply our algorithm of Sec. 16, greatly accelerated in the next section if $p(x)$ is given with its coefficients. In that case we immediately estimate the largest root radius $r_{1}$ to initialize Newton’s iterations of [115, 113, 102]. In this subsection we obtain such estimates for a black box polynomial.
@@ 24.5.1 Introductory comments
Given the coefficients of a polynomial $p$ of a degree $d$, we can readily compute its largest and smallest root radii $r_{1}$ and $r_{d}$, respectively (cf. [31, 39]). Next we compute a range $[\rho_{-},\rho_{+}]$ that brackets all root radii or, equivalently, compute an upper estimate for the largest root radius $r_{1}$ and a low estimate for the smallest root radius $r_{d}$ of a black box polynomial $p(x)$ of a degree $d$.
Its solution is required, e.g., in order to initialize algorithms for all $d$ roots by means of functional iterations.
Suppose that we have an algorithm that computes a disc $D(0,\rho_{-})$ containing no zeros of a black box polynomial $p(x)$; clearly, $\rho_{-}\leq r_{d}$. By applying this algorithm to the reverse polynomial $p_{\rm rev}(x)$ we obtain the upper bound $1/\rho_{-,{\rm rev}}$ on $r_{1}$.
@@
In the rest of this section $p(x)$ denotes a black box polynomial.
@@ 24.5.3 Deterministic solution
We can deterministically compute a desired disc $D(0,\rho_{-})$ at a higher computational cost, based on a soft e/i test.
Namely, apply 1-test to $d+1$ triples of parameters $c_{i}$, $\theta_{i}$, and $\rho_{i}$, for $i=0,1,\dots,d$, such that the discs $D(c_{i},\theta_{i}\rho_{i})$ have no overlap pairwise. Then, clearly, at most $d$ discs contain roots because $p(x)$ has only $d$ zeros, and so at least one of the $d+1$ discs contains no roots. Deterministic e/i tests of Secs. 7.2 and 7.3 involve evaluation of $p(x)$ at $22d+$ 2 points. We obtain a desired disc $D$ by applying these tests to at most $d+1$ discs and thus computing $p(x)$ at at most $2(d+1)^{2}$ points.
@@ 25 Fast approximation of the $d$ root radii with applications
The customary initialization recipes of Weierstrass’s and Ehrlich’s iterations involve the $d$ root radii at the origin and consequently require to use the coefficients of $p$. To initialize Ehrlich’s iterations MPSolve first applies Bini’s heuristic recipe of [6], reused in [9, 15], for the approximation of the $d$ root radii. Given such approximations MPSolve defines $k$ concentric annuli $A(c,\rho_{i},\rho_{i}^{\prime})$ and $k$ integers $m_{i}$ such that #$\#(A(c,\rho_{i},\rho_{i}^{\prime}))=m_{i}$ for $i=1,\dots,k$, $k\leq d$, and $\sum_{i=1}^{k}m_{i}=d$. Then MPSolve initializes Ehrlich’s iterations at $d$ points $z_{j}$, $j=1,2,\dots,d$, exactly $m_{i}$ of them being equally spaced on the circle $C(c,(\rho_{i}+\rho_{i}^{\prime})/2)$. This heuristic initialization recipe turns out to be highly efficient.
The algorithm supporting the next theorem also approximates the $d$ root radii. It appeared much earlier than Bini’s, is deterministic, formally supported, and faster than Bini’s. The algorithm has been devised and briefly analyzed in [106] for fast approximation of a single root radius and then readily extended in [74, Sec. 4] (cf. also [75, Sec. 5]) to approximation of all root radii of a polynomial. As in [6, 9, 15] it uses all $d$ coefficients of $p$. A variation of this algorithm in [47] supports the following theorem.
@@
[47, Prop. 3]. Given the coefficients of a polynomial $p=p(x)$, one can approximate all the $d$ root radii $|x_{1}|,\dots,|x_{d}|$ within the relative error bound $4d$ at a Boolean cost in $O(d\log(||p||_{\infty}))$.
@@
Combine Thm. 25.1 with estimates for the Boolean cost of performing these iterations implicit in the proof of [106, Cor. 14.3] and obtain
@@
The algorithm supporting the corollary and said to be Alg. 25.1a computes at a low cost $d$ suspect annuli $A(0,r_{j,-},r_{j,+})$ for $j=1,\dots,d$ (some may overlap), whose union $\mathbb{U}$ contains all $d$ roots. The algorithm was the point of departure for amazing progress in [44], which extends the ingenious algorithm of [64] for multipoint polynomial evaluation.
@@
(ii) The search for real roots can be narrowed to the intersection $\mathbb{I}$ of $\mathbb{U}$ with the real axis, which consists of $2m$ segments of small length (see Fig. 9). [98] has tested a real root-finder that approximates the roots in $\mathbb{I}$ by initiating Newton’s iterations at the centers of the segments. According to these tests the algorithm is highly efficient for the approximation of all well-isolated real roots. One can extend Newton’s iterations at the centers of the segments by applying also our root-radii algorithm of Sec. 16. By complementing the root radii approach with some advanced novel techniques, the paper [47] has noticeably accelerated the real root-finder [54], which is currently user’s choice.
(iii) Initialization based on Alg. 25.1a should improve global convergence of Ehrlich’s, Weierstrass’s, and other functional iterations for root-finding as well as of Kirrinnis’s factorization algorithm of [52].
(iv) Suppose that Alg. 25.1a has computed an annulus $A(0,\rho_{-},\rho_{+})$ containing a small number $u$ of the zeros of $p$ and isolated from its external zeros of $p$. Then we can extend the root-finding recipes of [106] from a disc to an annulus as follows. (a) Readily compute the first $u$ power sums of the zeros of $p$ lying in the discs $D(0,\rho_{+})$ and $D(0,\rho_{-})$ and then (b) their differences equal to the power sums of the zeros of $p$ lying in the annuls $A(0,\rho_{-},\rho_{+})$. (c) By solving the system of Newton’s identities, compute the coefficients of the factor $f$ of the $u$th degree that shares with $p$ its root set in the annulus (cf. [106, Sec. 12]). (d) Approximate the zeros of $f$, at a lower cost since $u$ is small.